A toy manufacturer needs a piece of plastic in the shape of a right triangle with the longer leg 5 cm more than the shorter leg and the hypotenuse 10 cm more than the shorter leg. How long should the sides of the triangle be
Answer by checkley79(3341)

A=B+5
C=B+10
A^2+B^2=C^2
(B+5)^2+B^2=(B+10)^2
B^2+10B+25+B^2=B^2+20B+100
2B^2-B^2+10B-20B+25-100=0
B^2-10B-75=0
(B-15)(B+5)=0
B-15=0
B=15 ANS. FOR THE SHORTER LEG.
A=15+5=20 ANS. FOR THE LONGER LEG.
C=15+10=25 ANS. FOR THE HYPOTENUSE.
PROOF:
20^2+15^2=25^2
400+225=625
625=625